OpenStudy (highschoolmom2010):

The radius of circle O is 18, and OC = 13. Find AB. Round to the nearest tenth, if necessary.

OpenStudy (highschoolmom2010):



OpenStudy (anonymous):

\[\huge BC^2 = OB^2 - OC^2 =18^2 - 13^2 \] \[\huge BC^2 = 18^2 - 13^2 =324 -169=155 \] \[\huge BC = \sqrt {155}\] Since AB = 2BC = \[2 \sqrt{155} = 2 \times 12.449 = 24.898 \approx 25\] @highschoolmom2010

OpenStudy (anonymous):